OpenStudy (sleepyjess):

Did I do this right? Finding the equation of a hyperbola, center not at (0, 0) Question below

OpenStudy (sleepyjess):

41. Center at (4, -1); focus at (7, -1); vertex at (6, -1) From that I figured out that I needed to use \(\dfrac{(x-h)^2}{a^2}-\dfrac{(y-k)^2}{b^2}=1\) Since 7-4 is 3, c = 3 6-4 is 2 a = 2 \(b^2\) = 9-4 \(b^2\) = 5 \(\dfrac{(x-4)^2}4 - \dfrac{(y+1)^2}9 = 1\)

OpenStudy (sleepyjess):

Whoops, that should be \(\dfrac{(y+1)^2}5\)

OpenStudy (phi):

yes, it looks correct

OpenStudy (sleepyjess):

Yay! Thank you :)

OpenStudy (mathmate):

Well done! http://prntscr.com/9ol6fa